A man is 71 and his daughter is 34. What age will the daughter be when she is half the age of her father? What age will the man be when he is 3 times the age of his daughter?
Answer by ankor@dixie-net.com(22740)

A man is 71 and his daughter is 34.
What age will the daughter be when she is half the age of her father?
Let y = no. of years for this to be true
y + 34 = .5(y+71)
y + 34 = .5y + 35.5
y - .5y = 35.5 - 34
.5y = 1.5
multiply both sides by 2
y = 3 yrs, Dad will be 74, daughter will be 37
:
What age will the man be when he is 3 times the age of his daughter?
Too late for this to happen in the future, in the past
3(34-y) = 71-y
102 - 3y = 71 - y
102 - 71 = -y + 3y
31 = 2y
y = 31/2
y = 15.5 years ago, Dad was 55.5, daughter was 18.5